# Embodied Image Quality Assessment for Robotic Intelligence

### v2

@@ V Experiment
V-A Experiment Setups
Oriented to the 100 episodes of the embodied task, 25 common image distortions are set up, and a total of 2,500 distorted images as well as the corresponding reference images are acquired. The train/val dataset is divided according to an 8:2 ratio. For each pair of images the overall scoring and the score of the two subtasks is annotated. Three metrics, SRCC, Kendall Rank Order Correlation Coefficient (KRCC) and Pearson Linear Correlation Coefficient (PLCC), are used to evaluate the consistency between objective and subjective quality scores. Where SRCC and KRCC represent predictive monotonicity and PLCC represents predictive accuracy.
For algorithm evaluation, we utilize 14 mainstream IQA methods for the comparison. These methods have achieved commendable results in terms of IQA methods based on the human visual system. In addition to comparing traditional PSNR and SSIM [1], the full reference IQA methods include AHIQ [24], PieAPP [25], CKDN [26], DISTS [27], TOPIQFR [28], IQT [29]. The no reference methods include CLIPIQA [30], DBCNN [31], TOPIQNR [28], MANIQA [32], TempQT [33], HyperIQA [34]. We trained 12 algorithms on the EPD dataset based on the Pyiqa [35] platform, following the data division described above. Instead of fixing the epoch during training, the models are trained until the loss converged relatively. We evaluated each algorithm on three evaluation metrics on three subdatasets. The hardware platform on which the experiments is 8$\times$×NVIDIA RTX3090 24G GPUs.
V-B Environment Setups
It is important to note that the EPD benchmark is different from the traditional paradigm. Instead of offline collection of various types of images, the EPD are oriented towards a real scenario with embodied AI. A robot needs to constantly perceive its surroundings as it completes its task. We capture images from a monocular camera mounted on the robot as it performs its task. Thus, the collected images are primarily for the robot and not for humans.
@@ V-B Environment Setups
Table I: Comparison of 14 IQA methods for BL (Baseline), FR (Full Reference), NR (No reference) respectively on EPD benchmarks. Blod represent the best results in BL. In FR and NR, the Best results are marked in Orange while the second best results are in Blue
Type Metric All Tasks Push Task Pick Task
@@
BL PSNR 0.1233 0.1356 0.0819 0.0811 0.0927 0.0539 0.0972 0.1138 0.0645
SSIM (TIP2004) [1] 0.0597 0.0635 0.0396 0.0633 0.0716 0.0417 0.0228 0.0275 0.0152
FR PieAPP (CVPR2018) [25] 0.3616 0.3853 0.2466 0.1727 0.2023 0.1165 0.1604 0.1802 0.1061
CKDN (ICCV2021) [26] 0.6971 0.6654 0.5062 0.2100 0.2186 0.1404 0.3372 0.3040 0.2266
IQT (CVPR2021) [29] 0.5435 0.5416 0.3814 0.3918 0.3706 0.2650 0.5613 0.5664 0.3920
AHIQ (CVPR2022) [24] 0.4199 0.4382 0.2888 0.2425 0.2707 0.1674 0.3061 0.3157 0.2074
DISTS (TPAMI2022) [27] 0.2113 0.2107 0.1428 0.1335 0.1608 0.0906 0.1135 0.1036 0.0985
TOPIQ-FR (TIP2024) [28] 0.1265 0.1232 0.0845 0.1615 0.1684 0.1113 0.1207 0.1185 0.0794
NR HyperIQA (CVPR2020) [34] 0.3212 0.3289 0.2212 0.3099 0.2981 0.2098 0.4055 0.3927 0.2733
DBCNN (TCSVT2021) [31] 0.1921 0.2133 0.1307 0.0962 0.1039 0.0643 0.1882 0.1855 0.1251
MANIQA (CVPR2022) [32] 0.5267 0.5603 0.3675 0.2475 0.2574 0.1661 0.5847 0.5859 0.4116
CLIPIQA (AAAI2023) [30] 0.1821 0.2160 0.1218 0.0750 0.0893 0.0503 0.1464 0.1497 0.0978
TempQT (TMM2024) [33] 0.2340 0.1730 0.1450 0.1040 0.0980 0.0560 0.2210 0.1400 0.1500
TOPIQ-NR (TIP2024) [28] 0.1253 0.1193 0.0828 0.0995 0.1043 0.0667 0.0846 0.0602 0.0568
@@
For the experiments, the ManiSkill benchmark [23] was chosen to set up the experimental environment, which is a development platform for reinforcement learning tasks and imitation learning with embodied intelligence based on the SAPIEN simulator [19]. The experiment is set up with two subtasks of pushing and picking a box, and the goal is to complete the specified tasks through reinforcement learning, i.e., pushing the box to the centre of the sign and picking the box. Each episodes represents one complete interaction of the robotic arm with the environment in different scenarios, and the experiment set 50 steps of actions for each episodes. Incidentally, reward values are given for each step.
@@
To explore the capability of the robot to task completion under different distorted images, the type and intensity of noise set in an episodes is consistent. The final task completion performance is measured by the average reward value. Thus, it can score the images captured by the robot while completing the embodied task, and the scores of several distorted images are shown in Fig. 5. Limited by the size and computational resources of the available reinforcement learning models, the image inputs captured by the robot are all 128$\times$128 pixels.
V-C Experiment Results and Discussion
@@
Table II: Comparison of MOS in different image distortion for 6 EAI participant. The D.1 to D.7 denote seven different types of image distortions, namely blus, colour distortions, compression, noise, brightness change, spatial distortions, sharpness and contrast. Bold data indicate the average best result
@@
Current state-of-the-art methods can only achieve performance below 0.7. In the FR method, CKDN [26] and IQT [29] are able to achieve SRCC and PLCC above 0.5. In the NR method only MANIQA [32] is able to achieve SRCC and PLCC above 0.5. Overall, the full-reference FR methods perform slightly better than the NR methods. Existing methods are essentially modelling the human visual system, while less attention is paid to the robotic visual system. It is interesting to note that CKDN [26] obtains state-of-the-art results on the EPD benchmark. CKDN uses the results of Image Restoration (IR) as a reference image for model training. This also gives us some inspiration that in embodied AI it is $only\ the\ machine\ that\ understands\ the\ machine\ better$.
@@
In addition, the performance of each method varies in different subtasks. The pick task has a higher performance than the push task, while the total task has the highest performance. EIQA still has significant optimisation space.
Table II shows the results of the average scores of distorted images from 6 different embodied AI participants. Each EAI assesses image quality differently in different image distortion environments. There are two participants each who scored high on noise and spatial distortion. One participant each scored the highest on compression, sharpness and contrast distortion. This also shows that there is variability in the image quality assessment by different EAI individuals.

### v3

@@ V Experiment
V-A Experiment and Environment Setups
Oriented to the 100 episodes of the embodied task, 25 common image distortions with 5 different levels of intensity are set up, and a total of 12,500 image pairs are acquired. The train/val dataset is divided according to an 8:2 ratio. For each pair of images the overall scoring and the score of the three subtasks (i.e. All Task, Push Task and Pick Task) is annotated by embodied robot. Three metrics, SRCC, Kendall Rank Order Correlation Coefficient (KRCC) and PLCC, are used to evaluate the consistency between objective and subjective quality scores. The SRCC and KRCC represent predictive monotonicity and PLCC represents predictive accuracy.
Figure 7: The different scenes that the benchmark can be supported. The EPD support push and pick two major types of tasks. A wide range of different objects and scenes can be flexibly changed. The first row demonstrates the reference undistorted image, while rows 2-5 demonstrate the corresponding distorted images. The experiment rendered various task scenarios from a third-person perspective. The model mainly takes the front view as input, such as 7-8 columns.
[…]
It is important to note that the EPD benchmark is different from the traditional paradigm. Instead of offline collection of various types of images, the EPD are oriented towards an embodied scene. A robot needs to constantly perceive its surroundings as it completes its task. We capture images from a monocular camera mounted on the robot as it performs its task. Moreover, the collected images are primarily for the robot and not for humans. The examples of diverse scenes are shown in Fig. 7.
@@ V-A Experiment and Environment Setups
Table I: Comparison of 16 IQA methods for BL (Baseline), FR (Full Reference), NR (No reference) respectively on EPD benchmarks. In FR and NR, the Best results are marked in Orange while the second best results are in Blue
Type Metric All Tasks Push Task Pick Task Params Architecture
@@
BL PSNR 0.1660 0.1107 0.1651 0.0928 0.0656 0.0969 0.1472 0.0978 0.1429 - Pixel-based
SSIM (TIP2004) [9] 0.2347 0.1572 0.2407 0.1392 0.0931 0.1386 0.2085 0.1390 0.2111 - Pixel-based
FR PieAPP (CVPR2018) [53] 0.2538 0.1707 0.1942 0.1863 0.1243 0.1414 0.2143 0.1438 0.2040 68.38 M CNN-based
LPIPS (CVPR2018) [56] 0.1886 0.1254 0.2062 0.1110 0.0736 0.1213 0.1915 0.1279 0.1898 14.72 M CNN-based
CKDN (ICCV2021) [54] 0.3671 0.2504 0.3435 0.2511 0.1688 0.1836 0.4310 0.2968 0.4118 13.38 M CNN-based
IQT (CVPR2021) [58] 0.5123 0.3568 0.5315 0.3654 0.2493 0.3838 0.5460 0.3839 0.5618 57.72 M Transformer
AHIQ (CVPR2022) [52] 0.3068 0.2080 0.3245 0.2553 0.1722 0.2647 0.3315 0.2254 0.3172 139.30 M Transformer
DISTS (TPAMI2022) [55] 0.0737 0.0495 0.0641 0.0625 0.0415 0.0685 0.0932 0.0627 0.0946 14.72 M CNN-based
TOPIQ-FR (TIP2024) [57] 0.2923 0.1974 0.3052 0.1425 0.0952 0.1471 0.2332 0.1560 0.2417 36.04 M CNN-based
NR HyperIQA (CVPR2020) [63] 0.4502 0.3081 0.4487 0.4604 0.3155 0.4686 0.4762 0.3324 0.4943 27.38 M CNN-based
DBCNN (TCSVT2021) [60] 0.2922 0.1971 0.2981 0.2851 0.1922 0.2894 0.3371 0.2295 0.3349 15.31 M CNN-based
MANIQA (CVPR2022) [61] 0.5526 0.3855 0.5716 0.4874 0.3368 0.5032 0.5666 0.4032 0.5956 135.75 M Transformer
CLIPIQA (AAAI2023) [59] 0.3471 0.2344 0.3222 0.2890 0.1941 0.2748 0.3322 0.2265 0.3152 102.02 M Transformer
TempQT (TMM2024) [62] 0.5500 0.3840 0.5620 0.5380 0.3760 0.5460 0.5470 0.3880 0.5780 87.45 M Transformer
TOPIQ-NR (TIP2024) [57] 0.2139 0.1433 0.2251 0.1081 0.0720 0.1227 0.1959 0.1308 0.2009 45.20 M CNN-based
QualiCLIP (CVPR2025) [64] 0.3728 0.2528 0.3669 0.3343 0.2261 0.3213 0.3669 0.2506 0.3713 102.12 M Transformer
MA-EIQA (our) 0.5755 0.4032 0.5836 0.4933 0.3408 0.5040 0.5738 0.4084 0.5971 48.83 M CNN-based
@@
For the experiments, the ManiSkill benchmark [48] is chosen to set up the experimental environment, which is a development platform for reinforcement learning tasks and imitation learning with embodied intelligence based on the SAPIEN simulator [44]. The experiment is set up with two subtasks of pushing and picking a box, and the goal is to complete the specified tasks through reinforcement learning, i.e., push the box to the center of the sign and pick up the box. Each episodes represents one complete interaction of the robotic arm with the environment in different scenarios, and the experiment set 50 steps of actions for each episodes. Incidentally, reward values are given for each step.
@@
To explore the capability of the robot to task completion under different distorted images, the type and intensity of distortion set in an episodes is consistent. The final task completion performance is measured by the average reward value. Furthermore, it can score the images captured by the robot while completing the embodied task. The example scores of several distorted images are shown in Fig. 8. Limited by the size and computational resources of the available reinforcement learning models, the image inputs captured by the robot are all $\times128$ pixels.
V-B Experiment Results on EPD
Table II: Score distribution for 25 common image distortions at level 5 distortion for embodied tasks. MEAN represents the average score for embodied tasks, and std represents the standard deviation of stability. The Best results are marked in Orange while the worst results are in Blue.
Distortion Type Distortion Level 1 Level 2 Level 3 Level 4 Level 5 Average
Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std
Blurs (D.1) Gaussian blur 2.8183 0.5226 2.7588 0.8119 2.6850 0.8089 2.5604 0.7954 2.2156 0.8319 2.6076 0.7541
Lens blur 2.7305 0.8097 2.7021 0.7760 2.4597 0.7622 2.2835 0.7404 2.1969 0.8697 2.4745 0.7916
Motion blur 2.8183 0.5226 2.7788 0.7892 2.7000 0.7819 2.7547 0.7417 2.5791 0.8173 2.7262 0.7305
Color distortions (D.2) Color diffusion 2.5685 0.8049 1.9424 0.8155 2.0131 0.7804 1.9942 0.7382 1.9553 0.7599 2.0947 0.7798
Color shift 2.6290 0.7970 2.4354 0.7373 2.4082 0.7590 2.4405 0.7676 2.4545 0.7972 2.4735 0.7716
Color quantization 2.7545 0.7629 2.7202 0.7623 2.6684 0.7791 2.6959 0.7908 2.4486 0.7456 2.6575 0.7681
HSV saturation 2.3168 0.8420 2.2403 0.8590 2.2804 0.7777 2.2837 0.8094 2.1825 0.8543 2.2608 0.8285
Lab saturation 2.8183 0.5226 2.2509 0.8297 2.1908 0.7793 2.1862 0.7977 2.2440 0.8146 2.3380 0.7488
Compression (D.3) JPEG2000 compression 2.7520 0.7035 2.6185 0.7749 2.6097 0.7316 2.5100 0.7677 2.1363 0.8032 2.5253 0.7562
JPEG compression 2.6127 0.8168 2.5582 0.7760 2.5606 0.7970 2.4506 0.7461 2.3529 0.8164 2.5070 0.7905
Noise (D.4) White noise 2.7048 0.8208 2.6860 0.7839 2.7622 0.7976 2.7136 0.7424 2.5121 0.8359 2.6758 0.7961
Color noise 2.7789 0.7491 2.8338 0.7379 2.7198 0.7771 2.6389 0.8048 2.5842 0.7263 2.7111 0.7590
Impulse noise 2.7791 0.7460 2.7432 0.7363 2.6892 0.7864 2.4918 0.8310 2.5504 0.7533 2.6507 0.7706
Multiplicative noise 2.6600 0.8230 2.7945 0.7941 2.7574 0.8312 2.6567 0.7343 2.5809 0.7816 2.6899 0.7928
Gaussian Denoise 2.6776 0.7739 2.5497 0.8320 2.4636 0.8194 2.3594 0.7262 2.3794 0.7772 2.4860 0.7857
Brightness change (D.5) Brighten 2.6550 0.8400 2.6327 0.8671 2.6563 0.7681 2.4700 0.8050 2.3525 0.8251 2.5533 0.8211
Darken 2.7048 0.7788 2.6715 0.7927 2.5773 0.8330 2.4169 0.8329 2.0651 0.8088 2.4871 0.8092
Mean shift 2.6778 0.7988 2.6805 0.8163 2.8183 0.5226 2.7004 0.7926 2.4519 0.8061 2.6658 0.7473
Spatial distortions (D.6) Jitter 2.7242 0.7540 2.7016 0.7544 2.6773 0.8158 2.6427 0.7981 2.6579 0.8887 2.6807 0.8022
Non-eccentricity patch 2.4136 0.7920 2.3908 0.7632 2.3191 0.8621 2.1849 0.8598 2.3457 0.7911 2.3308 0.8136
Pixelate 2.7576 0.7152 2.7173 0.8597 2.6423 0.7636 2.7744 0.8003 2.5421 0.7481 2.6867 0.7774
Quantization 2.7188 0.8034 2.7570 0.8052 2.6291 0.8201 2.6491 0.7754 2.5790 0.7973 2.6666 0.8003
Color block 2.2128 0.7638 2.2460 0.8536 2.0815 0.7723 2.1217 0.8157 2.1592 0.8185 2.1643 0.8048
Sharpness and contrast (D.7) High sharpen 2.6261 0.8346 2.5468 0.8290 2.4876 0.7981 2.4390 0.8280 2.3688 0.8509 2.4937 0.8281
Contrast change 2.5860 0.7816 2.5314 0.7457 2.6483 0.8699 2.6848 0.6908 2.4076 0.8465 2.5716 0.7869
@@ V-B Experiment Results on EPD
Traditional FR-based methods do not perform well because they differ from human assessment of images, whereas NR methods can extract more useful features from a robotic perspective.
Table III: Comparison of DMOS in different image distortion for 6 EAI participant, which normalised to between 0 and 5. The D.1 to D.7 denote seven different types of image distortions, namely blur, colour distortions, compression, noise, brightness change, spatial distortions, sharpness and contrast. Bold data indicate the average best result
@@
Current state-of-the-art methods can only achieve performance below 0.6. In the FR method, IQT [58] is able to achieve SRCC and PLCC above 0.5 on All Task dataset. In the NR method, MANIQA [61] and TempQT [62] are able to achieve SRCC and PLCC above 0.5. Models based on the Transformer architecture have higher average performance. Overall, the NR methods perform slightly better than the FR methods. Existing methods are essentially model the human visual system, while less attention is paid to the robotic visual system. The NR approaches can extract more useful features from a robotic perspective. This phenomenon further reveals a fundamental characteristic of embodied AI systems: onlythemachinethatunderstandsthemachinebetter.
@@
In addition, the performance of each method varies in different subtasks. The Pick task has a higher performance than the Push task, while the All task has the highest performance. Embodied IQA still has significant optimization space.
Table II shows the scores of 25 common image distortions in embodied tasks of level 5 normalized to (0, 5). The mean values in the table reflect the overall task performance, while the standard deviation std reflects the impact of image distortion on task robustness. To validate the influence of task types and participants on task results, six embodied intelligent robot subjects are selected, experiments are conducted as outlined in Table III. Just as different human participants have varying evaluations of images, so too do different embodied robots. Each EAI assesses image quality differently in different image distortion environments. Based on the results in the table, where findings are summarized as follows:
  • Different image distortions have varying degrees of impact on embodied tasks. Motion blur causes minimal disruption to embodied tasks, while color distortion causes maximum disruption.
  • Image distortion with minimal motion blur affects the robustness of embodied tasks, while color distortion reduces robustness.
  • The noise category of image distortion has the least impact on embodied tasks, while color distortion has the most severe impact on embodied tasks.
  • Embodied tasks have different sensitivities to changes in image distortion strength. JPEG2000 compression and Color diffusion are the two types of image distortion that are most sensitive to embodied tasks, while multiplicative noise is the least sensitive.
V-C Performance analysis of MA-EIQA
[…]
Different from offline network inference, embodied robots in industrial settings have limited computing resources. In the practice of deep learning deployment on edge-side embedded devices, CNN architectures benefit from long-term hardware adaptation and currently dominate traditional edge computing platforms. Transformer architectures, on the other hand, rely on matrix multiplication acceleration units in next-generation NPUs, and their industrial support varies. Image quality assessment should not consume excessive computational resources throughout the embodied task pipeline. Consequently, the adoption of a lightweight network architecture becomes imperative. Table. I compares the model parameters of our proposed method with those of the mainstream IQA methods. MA-EIQA is based on a CNN architecture with 48.83 million parameters. The number of parameters is reduced by 64.03 $\%$ compared to the Transformer-based MANIQA [61] and by 44.16 $\%$ compared to TempQT [62].
V-D Ablation Study
To systematically evaluate the effectiveness of each core component of the proposed MA-EIQA, we perform the ablation studies on EPD dataset. This experiment aims to verify the specific contributions of the multi-scale feature extraction module (MS) and the embodied attention module (EA) to the performance of model. All model variants used identical training strategies, hyperparameter configurations, and data splits to ensure a fair comparison. There the following four model for comparison:
Baseline: The backbone without the modules—the multi-scale feature encoder (MS) and the embodied attention module (EA).
Baseline+MS: The backbone with the modules—the multi-scale feature encoder (MS) but without the embodied attention module (EA).
Baseline+EA: The backbone with the embodied attention module (EA) but without the modules—the multi-scale feature encoder (MS).
MA-EIQA: Our comprehensive approach includes the modules—the multi-scale feature encoder (MS) and the embodied attention module (EA).
The four model are tested 10 times and the mean results are shown in Table. IV.
[…]
Table IV: Comparison of 14 IQA methods for BL (Baseline), FR (Full Reference), NR (No reference) respectively on EPD benchmarks. Blod represent the best results in BL. In FR and NR, the Best results are marked in Orange while the second best results are in Blue
Metric All Tasks Push Task Pick Task Average
[…]
Baseline 0.5138 0.3557 0.5246 0.4221 0.2892 0.4315 0.5303 0.3733 0.5464 0.4887 0.3394 0.5008
Baseline+MS 0.5370 0.3742 0.5521 0.4839 0.3338 0.4917 0.5412 0.3842 0.5580 0.5207 0.3641 0.5339
Baseline+EA 0.5281 0.3671 0.5347 0.4509 0.3091 0.4612 0.5436 0.3819 0.5585 0.5075 0.3527 0.5181
MA-EIQA 0.5755 0.4032 0.5836 0.4933 0.3408 0.5040 0.5738 0.4084 0.5971 0.5475 0.3841 0.5616
Figure 9: Cases of real-world experiment. The input for the robot is first-person view images with various types of distortion. The third-person perspective shows the results of the execution. We measure the error of task execution by calculating the Euclidean distance from the ground truth of human expert. The score is based on our MA-EIQA prediction score.
V-E Real-world Experiment
We conducted experiments on image distortion and embodied robot operation tasks in real-world scenes. We constructed a flexible experimental scene, including a display platform with arbitrary desktop background changes and variable environmental conditions. In the experiment, the UR5 robot is used as an embodied agent, and the coordinates of the end effector are controlled based on the ROS system. In the experiment, actual working conditions are simulated by changing the external environment and image distortion. Typical scenarios that affect embodied tasks include changes in illumination, object occlusion, lens coverage, and distortion from the camera sensor.
Fig. 9 demonstrates the first-person view image input under image distortion conditions and the results after the robotic arm performed the pick task. We use MA-EIQA to perform preference scores on input images. The ability to perform the task is evaluated based on the Euclidean distance between the results executed by the embodied agent and the position where the operation was successfully completed by an experienced human expert. As can be seen from the figure, compression and darkening have a greater impact on embodied tasks. Noise distortion and blocking have minor effects on embodied agents, whereas they severely affect human aesthetics. This also illustrates the difference between humans and robots in terms of image quality perception.